### v1

# U-MedSAM: Uncertainty-aware MedSAM for Medical Image Segmentation

Abstract
Medical Image Foundation Models have proven to be powerful tools for mask prediction across various datasets. However, accurately assessing the uncertainty of their predictions remains a significant challenge. To address this, we propose a new model, U-MedSAM, which integrates the MedSAM model with an uncertainty-aware loss function and the Sharpness-Aware Minimization (SharpMin) optimizer. The uncertainty-aware loss function automatically combines region-based, distribution-based, and pixel-based loss designs to enhance segmentation accuracy and robustness. SharpMin improves generalization by finding flat minima in the loss landscape, thereby reducing overfitting. Our method was evaluated in the CVPR24 MedSAM on Laptop challenge, where U-MedSAM demonstrated promising performance.
Keywords:
Medical Image Segmentation Segment Anything Model Uncertainty-aware Learning Sharpness-Aware Minimization
1 Introduction
Figure 1: Overview of the MedSAM [17].
Medical image segmentation is vital in clinical practice, enabling precise medicine [22], assessing therapeutic outcomes, and disease diagnosis by delineating organ boundaries and pathological regions, enhancing anatomical understanding, and abnormality detection [6]. Early segmentation models for medical images are typically designed for specific tasks with limited data. They might not handle the complex patterns and minute variations in medical images, which are often critical for clinical diagnosis and health science study [14].
Recently, foundation models, such as MedSAM [17] have been developed for medical image segmentation [10] [24] [25] (See Fig. 1), which are more strong, efficient, and applicable to different data modalities and situations. However, the uncertainty of the learning process associated with MedSAM on the diverse dataset has not been thoroughly investigated. By understanding this uncertainty, we can improve the robustness of MedSAM, making it more trustworthy for practical applications [20].
In this paper, we address the above challenge by proposing the U-MedSAM model, an extension of the MedSAM model [17], and improve its training by incorporating a novel uncertainty-aware hidden task learning. The uncertainty-aware framework enables the model to balance different aspects of the segmentation sub-tasks (e.g., boundary, pixel, region) adaptively, improving overall performance and robustness, and addressing issues like class imbalance[2]. Specifically, our newly introduced uncertainty-aware loss function integrates three components [8]: (1) Pixel-based loss: Measures differences at the pixel level (e.g., Mean Squared Error loss) [16]. (2) Region-based loss: Used for regions segmentation (e.g., Dice coefficient loss). (3) Distribution-based loss: Compares predicted and ground truth distributions (e.g., Cross-entropy loss and Shape-Distance loss). Finally, we utilize the Sharpness-Aware Minimization (SharpMin) optimizer [3, 14, 13, 11, 12] to improve the generalization of our U-MedSAM. We evaluate and compare our method against state-of-the-art MedSAM models, demonstrating superior performance. Our approach achieves optimal results by promoting sharper and more precise segmentation boundaries, thereby enhancing the accuracy and robustness.
In summary, we make the following key contributions:
  • We introduce an uncertainty-aware learning loss combining pixel-based loss (e.g., Mean Squared Error loss), region-based loss (e.g., Dice coefficient loss), and distribution-based loss (e.g., Cross-entropy loss and Shape-Distance loss). This design enables the model to conveniently incorporate any loss function for training the model jointly.
  • By employing auto-learnable weights [9] rather than fixed weights, our model can dynamically adjust the contribution of each loss component based on the uncertainty associated with each prediction. This approach allows the model to prioritize confident predictions and reduce the impact of ambiguous or noisy data.
  • By using the Sharpness-Aware Minimization (ShapMin) optimizer, our U-MedSAM finds parameter values that result in flat minima in the loss landscape. This technique improves the model’s ability to generalize and mitigates overfitting, a common issue in deep learning applications for medical imaging [18].
2 Method
This section will describe the entire workflow and implementation details. Fig. 2 shows an overview of our proposed method.
2.1 Preprocessing
We follow the preprocessing in LiteMedSAM implementation (baeline) [17].
Figure 2: Overview of our proposed U-MedSAM model: Building on the MedSAM architecture, we encode input images and introduce a novel uncertainty-aware learning method for automatic weight learning across multiple loss functions. The Sharpness-Aware Minimization (SharpMin) optimizer is employed, operating within a flattened loss landscape to enhance the generalization.
2.2 Proposed Method
2.2.1 Uncertainty-aware Auto-learning.
Uncertainty-aware learning enables the model to adapt its learning process based on detected uncertainties, enhancing resilience and precision by focusing on confident predictions and minimizing the influence of ambiguous ones [8]. This optimization of the training process across various datasets results in improved overall performance [26, 27].
By incorporating uncertainty into the loss computation, the model can dynamically assign weights to each loss component  [5]. This strategy effectively balances global and local accuracy while mitigating the impact of class imbalance. This approach allows the model to prioritize learning from reliable instances and reduce the impact of potentially erroneous or ambiguous data. Consequently, the model becomes more resilient to noisy or ambiguous data, leading to significantly improved segmentation performance [4, 21, 19].
The uncertainty-aware loss $\mathcal{L}$ if formulated as following:
$\mathcal{L}=\frac{1}{2\sigma_{1}^{2}}\mathcal{L}_{1}+\frac{1}{2\sigma_{2}^{2}}\mathcal{L}_{2}+\cdots+\frac{1}{2\sigma_{m}^{2}}\mathcal{L}_{m}+\cdots\log\sigma_{1}\sigma_{2}\cdots\sigma_{m},$ (1)
where $M$ is the number of individual loss components, $\mathcal{L}_{m}$ represents each individual loss component (such as DC, CE, MSE, and Shape-Distance loss), and $\sigma_{m}$ are learnable noise parameters. Larger noise levels indicate that the talk is not properly understood, necessitating greater learning weights. These learnable parameters are optimized during the training process to minimize the overall loss. The last term can constrain the noise to be increased too much. Specifically, we incorporate the following loss functions, see the survey [7] to find more losses.
  1. 1.
    Mean Squared Error (MSE) loss: This pixel-based loss ensures accurate classification of individual pixels by calculating the squared difference between each pixel’s predicted value and the true label to quantify the model’s prediction accuracy [7].
  2. 2.
    Dice Coefficient (DC) loss: This region-based metric emphasizes the overlap between predicted and ground truth areas, preserving the accuracy of the shape and boundaries of segmented regions [7].
  3. 3.
    Cross-Entropy (CE) loss: This distribution-based loss ensures accurate categorization of individual pixels, improving classification precision [7].
  4. 4.
    Shape-Distance (SD) loss: This distribution-based loss enhances shape features in image segmentation, strengthens boundary alignment, and ensures that the model pays closer attention to the geometric and structural information of the targets during the training process [1].
2.2.2 Sharpness-Aware Minimization.
Our approach employs Sharpness aware Minimization (SharpMin) optimization [3, 20, 15, 11, 14, 12] to enhance the generalization ability of U-MedSAM. By flattening the loss landscape, Sharpness-Aware optimization improves the model’s generalization potential. Traditional optimization techniques target the lowest points in the loss landscape, which can be steep and result in poor generalization on new data. In contrast, Sharpness-Aware Minimization identifies smoother minima, regions in the parameter space where the model’s performance remains consistent and less susceptible to disturbances.
2.3 Post-processing
No post-processing techniques were applied, and no strategies were implemented to enhance inference speed.
3 Experiments
3.1 Dataset and evaluation measures
We only used the challenge dataset for model development. The evaluation metrics consist of two accuracy measures: the Dice Similarity Coefficient (DSC) and Normalized Surface Dice (NSD), along with an efficiency measure of running time. Together, these metrics contribute to the overall ranking computation.
3.2 Implementation details
3.2.1 Environment settings
The development environments and requirements are presented in Table 1.
Stage 1. Code and models are available at: https://github.com/liangzw599/Co-developed-by-LiteMedSAM.
Stage 2 (Post challenge analysis). Code and models are available at: https://github.com/liangzw599/Co-developed-by-LiteMedSAM.
Table 1: Development environments and requirements.
System Ubuntu 20.04.1 LTS
CPU Intel(R) Xeon(R) CPU E5-2690 v4 @ 2.60GHz
RAM 125GB; 1.03MT///s
GPU (number and type) Four NVIDIA Titan Xp
CUDA version 11.4
Programming language Python 3.10.14
Deep learning framework torch 2.1.2, torchvision 0.16.2
Table 2: Training protocols.
Pre-trained Model SAM [10] MedSAM [17]
Batch size 2
Patch size 256$\times$256$\times$3
Total epochs 135
Optimizer AdamW
Initial learning rate (lr) 0.00005
Lr decay schedule ReduceLROnPlateau (mode = min,factor = 0.9,
patience = 5, cooldown = 0)
Training time 370 hours
Loss function Dice loss, BCE loss, SD loss and IoU loss
Number of model parameters N/A - not tracked
Number of flops N/A - not tracked
3.2.2 Training protocols
We follow the training protocols in LiteMedSAM implementation (baseline) [17].
4 Results and discussion
Table 3: Quantitative evaluation results.
Target LiteMedSAM Only SD-loss No SharpMin U-MedSAM
DSC(%) NSD(%) DSC(%) NSD(%) DSC(%) NSD (%) DSC(%) NSD (%)
CT 89.10 91.03 87.20 89.63 89.61 91.66 89.89 91.69
MR 83.28 86.10 79.27 82.01 81.68 83.81 82.79 84.61
PET 55.10 29.12 65.23 47.23 66.34 50.17 66.63 48.71
US 94.77 96.81 83.40 88.63 84.30 89.27 85.23 90.38
X-Ray 75.83 80.39 78.59 84.31 83.22 88.94 85.28 90.14
Dermotology 92.47 93.85 93.83 95.32 93.89 95.34 94.10 95.58
Endoscopy 96.04 98.11 93.60 96.24 95.65 97.93 96.24 98.33
Fundus 94.81 96.41 95.45 97.05 94.87 96.50 95.66 97.26
Microscopy 61.63 65.38 78.20 84.59 79.92 86.64 79.05 85.33
Average 82.56 81.91 83.86 85.00 85.48 86.64 86.10 86.89
Table 4: Quantitative evaluation of segmentation efficiency in terms of running time (s). Note: The inference process cannot use GPU.
Case ID Size Num. Objects Baseline Proposed
3DBox_CT_0566 (287, 512, 512) 6 376.4 286.3
3DBox_CT_0888 (237, 512, 512) 6 100.5 88.7
3DBox_CT_0860 (246, 512, 512) 1 17.7 18.5
3DBox_MR_0621 (115, 400, 400) 6 157.1 146.7
3DBox_MR_0121 (64, 290, 320) 6 99.9 73.5
3DBox_MR_0179 (84, 512, 512) 1 17.1 9.8
3DBox_PET_0001 (264, 200, 200) 1 12.1 13.2
2DBox_US_0525 (256, 256, 3) 1 6.3 5.4
2DBox_X-Ray_0053 (320, 640, 3) 34 7.3 6.5
2DBox_Dermoscopy_0003 (3024, 4032, 3) 1 6.5 3.2
2DBox_Endoscopy_0086 (480, 560, 3) 1 6.1 7.3
2DBox_Fundus_0003 (2048, 2048, 3) 1 6.1 7.5
2DBox_Microscope_0008 (1536, 2040, 3) 19 6.8 6.4
2DBox_Microscope_0016 (1920, 2560, 3) 241 19.1 15.9
4.1 Quantitative results on validation set
Table 3 summarizes these results. The results demonstrate that incorporating uncertainty-aware loss and the SharpMin optimization indeed improves segmentation accuracy.
Training using only the SD loss yields 83.86% DSC. In comparison, training using the uncertainty-aware loss (which incorporates the DC loss, CE loss, and focused loss), enhances DSC to 85.48%. This result shows the capability of the uncertainty-aware loss to increase the robustness and accuracy of models by prioritizing confident predictions and minimizing the impact of uncertain ones. It effectively addresses the limitations of using a single loss function by balancing various aspects of the segmentation task.
Incorporating the SharpMin optimization into the uncertainty-aware loss yields 86.10% DSC, the highest among the examined approaches. SharpMin improves segmentation accuracy by generating clearer and more accurate borders, guiding the optimization process toward flat minima in the loss landscape. This leads to better generalization across diverse data samples. The increase in DSC from 85.48% to 86.10% highlights the effectiveness of SharpMin in enhancing segmentation results and more precisely identifying heart components.
4.2 Qualitative results on validation set
We show some examples with good segmentation results in Fig. 3 and examples with bad segmentation results in Fig. 4.
Figure 3: Qualitative results of well-segmented from various modalities.
4.3 Segmentation efficiency results on validation set
As shown in Table 4, we compared the running speed of the proposed method to the LiteMedSAM (baseline) [17].
4.4 Results on final testing set
This is a placeholder. We will announce the testing results for Post-challenge analysis.
4.5 Limitation and future work
Despite achieving promising dice scores, we did not optimize the model’s speed, encountering limitations in this area.
In future work, we will explore various model compression techniques to address these speed constraints. Techniques such as quantization and tensor compression will be investigated to enhance the model’s efficiency. Additionally, we will examine other advanced methods and approaches to further improve the performance and speed of the model. Our goal is to balance high accuracy with faster processing times, ensuring the model’s effectiveness and efficiency for practical applications. By integrating these enhancements, we aim to develop a more robust and versatile model capable of handling real-world scenarios with greater ease and reliability.
Figure 4: Qualitative results of challenging examples from various modalities.
5 Conclusion
We present U-MedSAM, an innovative model engineered for robust medical image segmentation. This model combines the MedSAM architecture with an uncertainty-aware learning framework to dynamically adjust the contribution of multiple loss functions. By employing the SharpMin optimizer, U-MedSAM is steered towards flat minima in the loss landscape, enhancing its resilience and improving generalization. Comparative analysis with top-performing models highlights U-MedSAM’s superior accuracy and robustness in segmentation performance.
5.0.1 Acknowledgements
We thank all the data owners for making the medical images publicly available and CodaLab [23] for hosting the challenge platform.
Xin Wang is supported by the University at Albany Start-up Grant.